# Investigating generalization capabilities of neural networks by means of loss landscapes and Hessian analysis

### v1

@@ 2.4 Loss landscapes of popular NNs
In this Section we examine loss landscapes of popular networks to investigate the applicability of the existing visualization methods. These are obtained for randomly initialized and ImageNet-1K pre-trained NNs. All models and their pre-trained weights are taken from PyTorch library [19] with the exception of VIT [20].
@@ 3.2 Typical HESD types
Figure 7: HESD plots of neural networks trained to 70-80% accuracy: (left) SqueezeNet 1.1 pre-trained on ImageNet-1K, (center) MobileNetV2 pre-trained on ImageNet-1K, and (right) LeNet trained on MNIST.
Such HESD plots have been observed for a large class of convolutional and fully-connected NNs. Similar HESD plots were also observed for Generative Pretrained Transformers (GPTs) [29]. It was suggested that typical HESD structures can be attributed to block structure of NNs and the same loss function (cross-entropy) used for all classifiers, which also holds true for our results.
@@
Lastly, whereas HESD of randomly initialized VIT behaves as expected, HESD of pre-trained VIT shown in Figure 9 is different from the previously discussed ones. It is characterized by a large negative eigenvalue section, even for VITs that have relatively high training accuracy and low loss. This is also peculiar since VIT loss landscape shown in Figure 3 does not show any large negative curvature. Whereas this observation is very interesting, it requires additional study and the detailed analysis of VIT behavior will be conducted later.
Figure 9: HESD plot of (left) randomly initialized and (right) ImageNet-1K pre-trained VIT.
@@ 3.3 Hessian analysis criteria
Table 2 also shows that re decreases for pre-trained models, too. However, its variation range is extremely large making it less convenient to use as a criterion compared to KH. It also sometimes does not reflect the changes in accuracy correctly, as will be shown in Section 4. On the other hand, it can help to capture abnormalities like the behavior of randomly initialized ResNet18 in train mode, for which re sky-rockets to 451. It should also be noted that the strange behavior of VIT HESD discussed previously results in all criteria giving opposite results for VIT compared to CNNs.
@@ 4.4 Generalization capability of popular NNs
VIT ImageNet 87.5 0.99 0.96 6.3
VIT NI 89.1 0.95 0.82 0.77
Table 4 shows that KH05 behaves as has been previously discussed in this study, increasing when generalization accuracy drops for CNNs. As mentioned earlier, KH1 does not always have this property, as AlexNet results show. In many cases re also increases when accuracy decreases, though these changes vary significantly for different experiments, and re behavior does not always reflect that of the accuracy. This makes KH05 the most suitable Hessian criterion for NN generalization capability assessment.
However, VIT is once again an exception. It should also be noted that NI accuracy being higher than ImageNet accuracy for VIT can be explained by the fact that the class set in NI is very unambiguous, resulting in a simpler tasks compared to the analysis of random ImageNet batches.
@@ 5 Conclusions
This paper discusses the applicability of loss landscape and Hessian analyses to the study of generalization capabilities of NNs while focusing on different aspects of the methodology aiming to propose new analysis criteria. The results were obtained using a Loss Landscape Analysis library developed as part of this study. Different approaches to loss landscape evaluation were discussed showing that conventional methods might fail in specific circumstances which raises the demand for an improved normalization method. This conclusion was made after a comprehensive analysis of loss landscapes of conventional NNs in various regimes was conducted. It was shown that using Hessian axes can sometimes mitigate this issue, and the approach for choosing Hessian axes was proposed. Typical Hessian spectra were also discussed and several criteria for their analysis were proposed, while highlighting the differences between CNNs and VIT. Generalization experiments conducted using LeNet and ResNet20 on MNIST-SVHN and cifar-cinic dataset pairs showed that the proposed criteria can be used as an estimate of generalization accuracy, which is computationally efficient for large datasets.

### v2

@@ 2.4 Loss landscapes of popular NNs
In this Section we examine loss landscapes of popular networks to investigate the applicability of the existing visualization methods. These are obtained for randomly initialized and ImageNet-1K pre-trained NNs. All models and their pre-trained weights are taken from PyTorch library [19] with the exception of VIT that was trained as part of this study.
@@ 3.2 Typical HESD types
Figure 7: HESD plots of neural networks trained to 70-80% accuracy: (left) SqueezeNet 1.1, (center) MobileNetV2, and (right) VIT trained on ImageNet-1K.
Such HESD plots have been observed for a large class of convolutional and fully-connected NNs. Similar HESD plots were also observed for VITs and Generative Pretrained Transformers (GPTs) [28]. It was suggested that typical HESD structures can be attributed to block structure of NNs and the same loss function (cross-entropy) used for all classifiers, which also holds true for our results.
@@
Lastly, HESD of pre-trained VIT [29] shown in Figure 9 is different from the previously discussed ones. It is characterized by a large negative eigenvalue section, even that this VIT has a relatively high training accuracy and low loss. It is likely to be related to the differences in training of this specific VIT [30] and other models discussed in this study. It requires additional study and the detailed analysis of this VIT’s behavior will be conducted later.
Figure 9: HESD plot of (left) randomly initialized and (right) pre-trained VIT from [29, 30].
@@ 3.3 Hessian analysis criteria
Table 2 also shows that re decreases for pre-trained models, too. However, its variation range is extremely large making it less convenient to use as a criterion compared to KH. It also sometimes does not reflect the changes in accuracy correctly, as will be shown in Section 4. On the other hand, it can help to capture abnormalities like the behavior of randomly initialized ResNet18 in train mode, for which re sky-rockets to 451.
@@ 4.4 Generalization capability of popular NNs
VIT ImageNet 98.5 0.86 0.47 0.04
VIT NI 4.9 0.99 0.93 0.49
Table 4 shows that KH05 behaves as has been previously discussed in this study, increasing when generalization accuracy drops for CNNs. As mentioned earlier, KH1 does not always have this property, as AlexNet results show. In many cases re also increases when accuracy decreases, though these changes vary significantly for different experiments, and re behavior does not always reflect that of the accuracy. This makes KH05 the most suitable Hessian criterion for NN generalization capability assessment. It should be noted that the studied VIT has generalized poorly, which is reflected in hessian criteria. The reasons for this will be investigated in detail in the future.
@@ 5 Conclusions
This paper discusses the applicability of loss landscape and Hessian analyses to the study of generalization capabilities of NNs while focusing on different aspects of the methodology aiming to propose new analysis criteria. The results were obtained using a Loss Landscape Analysis library developed as part of this study. Different approaches to loss landscape evaluation were discussed showing that conventional methods might fail in specific circumstances which raises the demand for an improved normalization method. This conclusion was made after a comprehensive analysis of loss landscapes of conventional NNs in various regimes was conducted. It was shown that using Hessian axes can sometimes mitigate this issue, and the approach for choosing Hessian axes was proposed. Typical Hessian spectra were also discussed and several criteria for their analysis were proposed. Generalization experiments conducted using LeNet and ResNet20 on MNIST-SVHN and cifar-cinic dataset pairs showed that the proposed criteria can be used as an estimate of generalization accuracy, which is computationally efficient for large datasets.